# ChemKANs for Combustion Chemistry Modeling and Acceleration

### v1

@@ 3.1.2 Neural scaling with noise-free data
Neural scaling is an effective method of measuring parameter efficiencies in neural networks. Here different parameter sizes were investigated by changing the number of nodes in the hidden layers. Training and test losses were evaluated for data without noise to isolate the underlying expressive capabilities of each technique. Figure 4 reveals neural scaling at orders of 1.0 and 0.6 for the training and testing metrics of ChemKANs (where the order of neural scaling is defined as the power to which the loss decreases with respect to the number of parameters). While we might expect values up to 4 based on prior KAN [23] and KAN-ODE [27] studies, we remark here that the low-order neural scaling appears to indicate relatively saturated training of the ChemKAN with errors already around with just 72 parameters, rather than poor convergence which might otherwise have been indicated if low convergence rates were coupled with poor loss metrics.
@@
To further probe the nuances of the ChemKAN’s convergence efficiency, we compare its results with DeepONet. The ChemKAN iterates roughly an order of magnitude slower but converges in fewer epochs (as originally discussed in [27]). Thus, to facilitate fair comparison the ChemKAN was trained only for 5,000 epochs, while the DeepONet was trained for 50,000 epochs (10 times more than ChemKAN). Two key distinctions between ChemKAN and DeepONet are as follows. First, as discussed earlier, the extremely sparse ChemKANs (toward the left half of Fig. 4(A)) are able to reach remarkably low error even with just 78 parameters, while the DeepONets see significantly worse performance at sparse parameterizations (i.e., the left half of Fig. 4(A)). Secondly, we note that the DeepONet sees significantly higher order neural convergence in the training results, allowing it to surpass the training performance of the ChemKAN at above 200 parameters, with remarkably strong training accuracy for the largest, 456-parameter DeepONet studied here. Linear fits with slopes are shown in Fig. 4(A) to illustrate this point, where the last ChemKAN and last two DeepONet points are excluded as they begin to plateau in training loss. From this observation, a large-enough DeepONet seems to outperform ChemKAN when looking only at the training losses.
@@
To further contextualize the structural efficiency, we plot results for the testing error of the same neural convergence runs in Fig. 4(B). Here, we notice a significant departure from the training convergence rates in Fig. 4(A). Unlike the neural convergence for training data, the ChemKAN is seen to outperform the DeepONet at all sizes, with the largest ChemKAN leveling out and retaining nearly the same testing performance as the second-largest ChemKAN. The DeepONets, while enjoying a faster neural convergence rate below 308 parameters, notably fail to plateau and instead appear to diverge at higher parameter counts. When compared against the training results in Fig. 4(A), we observe two distinct modes of training saturation. Saturation, or the point where the linear fit no longer holds, appears to occur for the two largest DeepONets. For the ChemKAN, we might either interpret the entire profile to be saturated, or highlight the single largest network as the saturation point. Regardless, what we observe in these high-parameter networks is high robustness in the ChemKAN to overfitting (with a flat testing loss plateau), compared to the significant overfitting and divergence seen in the DeepONet past 308 parameters (i.e., the last two points).
@@
It is unsurprising that a standard deep learning technique begins to overfit a small dataset when given a large number of parameters. The DeepONet overfitting past saturation leads to further decreases in training loss accompanied with significant increases in testing loss. What we do find surprising, however, is the ChemKAN’s apparent resilience to overfitting, even with similar parameter counts. In this context, its original failure to reach the same training performance as the DeepONet appears to be a strong point of the method rather than a drawback, as it reaches a minimum value for both training and testing and then remains robust to superfluously added parameters, while the DeepONet clearly requires additional care to avoid overfitting. This pilot neural convergence test suggests that ChemKANs are robust to overfitting, and motivates further study of their capability in a second, more realistic model inference scenario.
3.1.3 Benefits of latent-dynamics modeling: noisy inference comparison
@@
To gain deeper understanding of how ChemKAN outperforms DeepONet, we further test training behaviors of ChemKAN and DeepONet with noisy data. Early stopping is carried out for both network structures at 10,000 epochs, to enable fair comparison and limit overfitting in the DeepONet cases. In all cases, a DeepONet with 308 parameters was used to compare against a ChemKAN with 156 parameters. The 308-parameter DeepONet was chosen based on the results of Fig. 4, where this was seen to be the largest DeepONet before the test loss began overfitting. The 156-parameter ChemKAN, meanwhile, was chosen to roughly match the DeepONet’s training and testing performance at zero noise. In other words, we chose the best-performing DeepONet size possible based on the preliminary neural convergence study, and then sized the ChemKAN according to the zero-noise performance of both networks. In more detail, the DeepONet had a three-layer branch network with eight nodes per layer and a two-layer trunk network with seven nodes in the first layer and eight in the second layer, with a final output layer converting these two eight-dimensional layers to the six-dimensional solution vector. The ChemKAN, meanwhile, had a single hidden layer with four nodes, two of which included multiplication operators () as per the recommendation of [33], and three gridpoints per activation.
@@
Figure 5(A) shows average training results after epochs across the 20 training and 10 testing cases at varied noise levels from to . As discussed in Sec. 3.1.2, the training MSE with 0% noise shows that the 308-parameter DeepONet slightly beats the 156-parameter ChemKAN in training performance, and is slightly worse in terms of reconstructing the unseen testing data.
@@
As increasing noise is added to the system, we see in the standard training and testing metrics of Fig. 5(A) that both networks unsurprisingly see increases in training and testing errors. Roughly, the increase in MSE error in the training losses scales with the square of the noise, as follows from Eq. 19. For example, the ChemKAN sees an increase in MSE between 0% noise and 1% noise of , where the second degree scaling suggests that a 25 larger increase of might be expected between 0% noise and 5% noise. This is indeed observed, with an increase in this latter case of . In other words, the increase in training error for both networks as noise is added can be attributed to the effect of the noise itself on the loss function (Eq. 19), and does not appear to indicate any problems with the two networks’ capabilities to fit the increasingly noisy training data. For a direct comparison, The ChemKAN retains lower testing error throughout all tested noise values, and at noise and above is actually able to reach a lower training error than the DeepOnet. Looking at the big picture, however, results in these two metrics remain within a factor of two of each other at all noise levels, indicating largely similar performance.
@@
Upon further evaluation, we found that the training and testing MSEs evaluated on noisy datasets do not fully capture the effects of noise on useful model predictions, as overfitting can occur not only to the training conditions but also to the noise present in the data. To more effectively compare these frameworks, we introduce a noise-free MSE metric (similar to Ref. [30]),
@@
which differs from the test MSE in Eq. 19. This noise-free metric serves to quantify the capability of the two modeling approaches to accurately extract the true underlying model from noisy data, rather than overfit the noise or otherwise fail to deliver a useful model. With this metric, much more significant performance shifts can be seen in the noise-free testing values in Fig. 5(A), which are the true indicators of the two approaches’ capabilities to extract hidden underlying signals from noisy data. While the added noise clearly increases the training and testing error when evaluated on the noisy data itself for both model types, the impact of the added noise on the noise-free testing error for the ChemKAN is relatively small (a roughly increase from to noise), suggesting that it is able to extract the true underlying behavior well from the noisy datasets. We reiterate the significance of this result in the context of Ref. [30], where standard KANs were demonstrated to fail when faced with small added noise (while here the ChemKAN retains good performance). In contrast, the DeepONet sees a 5 increase in noise-free testing MSE from to noise, with its final noise-free testing loss 4.4 larger than that of the ChemKAN.
@@
The loss profiles in Figs. 5(B) provide further insight on the training dynamics that lead to this significant discrepancy in noise-free testing reconstructions. In the noise training cycle of Fig. 5(B), we see fairly standard behavior across the training and noise-free testing traces for the DeepONet and ChemKAN, with all quantities steadily decreasing for the entire duration of training. This is the expected result, as we have sized the DeepONet based on Fig. 4 to avoid any overfitting in the noise-free case. As we increase the amount of noise to , the training loss values are heavily penalized (due to the noisy data used in the computation of Eq. 19), while the noise-free testing values are slightly penalized but remain comparatively strong, indicating that both approaches are at least to a certain extent able to extract the true underlying model from the noisy data. While the training loss drops quickly and then plateaus in all four subplots, we see in the 7% noise case of Fig. 5(B) that the DeepONet noise-free testing loss dynamics begin to suffer, with a minimum value near 5,000 epochs and a slight upward trend toward later epochs, likely due to overfitting. In the 15% noise case of Fig. 5(B) this issue is further exacerbated, with an early minimum near 1,000 epochs followed by significant overfitting to the noisy data for the remainder of the training profile. The ChemKAN in both cases remarkably continues to drop its noise-free testing loss even while the DeepONet is overfitting, with late-epoch dynamics showing plateaued minimum values rather than the overfitting seen in the DeepONet. This echos the behavior seen in Fig. 4, where the ChemKAN did not overfit and instead simply plateaued at its minimum training and testing errors.
@@
Reconstructed training data profiles are shown for the 15% noise case in Fig. 6. While the DeepONet and ChemKAN are both roughly able to find the unseen, noise-masked ground truth profile, a close inspection reveals not only better ChemKAN fits, but also notably jagged profiles from the DeepONet as it attempts to overfit the noise present in the training data. These results suggest that the dynamical system exploitation inherent to the ODE-based framework of ChemKANs (and KAN-ODEs in general) are able to mitigate or even entirely resolve the issues observed in Ref. [30] regarding noisy data with KAN networks.
@@
In summary, we have demonstrated ChemKANs as a promising tool for model discovery in temperature-dependent chemical kinetic systems, especially with realistically noisy datasets. They show promise not only on the scientific side of the problem, where they were demonstrated here to have significant capability compared to a standard tool in discovering models hidden under synthetic noise, but also on the machine learning side of the problem, where we have demonstrated that the neural ODE implementation of KAN-ODEs and ChemKANs helps them to overcome the noisy data limitation recently shown in vanilla KAN networks [30].
3.2 Hydrogen combustion acceleration from homogeneous reactor data
@@
In our second case study, we investigate the use of ChemKANs as a reduced-order solver acceleration framework for known chemical models. We hypothesize based on previous high-order neural scaling results [27, 23] as well as the strong performance at low parameter counts seen in Fig. 4 that the use of a KAN structure instead of an MLP will allow for similar accuracy in network predictions with fewer parameters and lower cost. We additionally aim to evaluate the novel network architecture of the ChemKAN, and whether its physics-based, single-network structure is able to model behavior that required MLPs in the Neural ODE framework of Ref. [17]. To reduce computational cost to the furthest extent possible, we use a single hidden layer of just three nodes for the core kinetic network (Eq. 13). This three-node hidden layer works as a latent representation of the system’s dynamics, compressing the information from 10 thermochemical states in the input and 29 reactions in the kinetic model. Leveraging our knowledge of the functional form of the governing equations (Eqs. 1 and 2) and their strong multiplicative behavior, we use here, defining all three hidden nodes using the multiplication operator.
@@
We begin in Fig. 7 with a demonstration of the ChemKAN homogeneous reactor reconstructions for one training case (top row, and K) and the unseen testing case (bottom row, and K), after the two-stage training process outlined in Sec. 2.4.2. These reconstructions were generated entirely by the ChemKAN given only the initial conditions. Overall the learned model successfully predicted the temperature and mass fractions in both cases, with no notable deterioration in the testing case (as expected from the strong testing results and robustness to overfitting observed in the previous biodiesel investigation). We further emphasize that the ChemKAN was largely able to capture the behavior of the low-concentration and highly reactive species , , and (Fig. 7(C1-C2)) that were neglected in ChemNODE [17].
@@
A broader comparison is shown in Fig. 8(A), where the MSE (in the normalized u units) is plotted across not only the initial set of 35 training and 1 testing initial conditions, but a wider set of 441 total initial conditions (406 of which were unseen during training) at a finer resolution in the same range. The low-temperature initial conditions see near-perfect reconstructions, as the ignition delays there are larger than the studied time window, leading to smooth gradients and easily trainable, near-isothermal behavior in those cases. In the remainder of the domain, strong performance is seen at all training conditions, and additionally at the single testing condition plotted in Fig. 7. In terms of generalization to intermediate temperatures and equivalence ratios, the ChemKAN performed very well throughout the vast majority of the domain, with many testing points even surpassing the accuracy of nearby training points at and above 1050K. We notice that toward the slower-igniting cases, however, the ChemKAN struggled more with generalization. At 1000K, for example, all six training points saw strong MSE values in the range, although a few of the intermediate equivalence ratios suffered. At 987.5K, however (one tick below 1000K), all testing points saw poorer performance in the range. We can conclude from these results that the ChemKAN retains a strong capability to generalize in the more challenging H2-air combustion case (as was originally reported in the biodiesel modeling case), but with practical limits. At higher temperatures it was remarkably even seen to outperform training results in broad swaths of unseen testing conditions, but in the more initiation-sensitive low-temperature cases, it was only able to perform well locally near the training points, indicating that denser training data is needed to optimize performance in those regions.
@@
We finally plot the actual and KAN-ODE predicted ignition delays in Fig. 8(B), for the 30 studied cases that ignited (the lowest temperature cases did not see ignition given the time span of 0.6 s). Ignition here is defined as the point of maximum temperature rise rate, as in Ref. [17]. Accuracy is strong across the board, even in the testing case.
@@
This collection of results shows that the ChemKAN structure was able to accurately learn the dynamics of all nine species and temperature scalars across the same set of initial reactor conditions as was studied using traditional MLP-based Neural ODEs in ChemNODE [17]. Compared to the six species plus temperature scalars learned there via seven unique MLP networks with 91 parameters each according to standard MLP parameterizations (637 total parameters), the current ChemKAN was able to learn the complete set of thermochemical scalars (nine species plus temperature) using a single, 344-parameter network. While training took place in two stages to decouple the kinetic and thermodynamic behavior and facilitate convergence, the final network remains a single cohesive structure with shared information across all nodes, eliminating the redundancies in repeated yet isolated 91-parameter MLPs.
@@
Finally, regarding computational efficiency, we report that the average time to solve all 36 homogeneous reactor conditions 100 times each (3600 total solves) in the Arrhenius.jl combustion solver package [50] was 2 faster on average when switching from the detailed chemistry to the reduced ChemKAN framework. By learning the relationship between the current thermochemical state and the chemical source terms, the ChemKAN is capable not only of predicting ignition delay times and homogeneous reactor solution profiles 2 faster than the detailed model, but also of generalizing to other simulation conditions including simple laminar flames and complex 2-D and 3-D turbulent combustion conditions. Such downstream uses of similar surrogate machine learning models were discussed and tested in Refs. [16] and [17], where a 2 speedup in the chemical solver (which is often the most computationally expensive component in a reacting flow simulation) implies the potential for substantial acceleration unlocked by ChemKANs while retaining the full-sized solution state vector. A summarized comparison of ChemKAN and ChemNODE is provided in Table 1.
| ChemNODE | ChemKAN | |
| No. Networks | 7 | 1 |
| Parameters | 637 | 344 |
| Species modeled | H2, O2, H2O, N2, O, OH | H2, O2, H2O, N2, O, OH, H, HO2, H2O2 |
| Speed-up vs. full chemistry | 2.3x | 2.0x |
3.3 Current limitations and areas for further research

### v2

@@ III.1.2 Neural scaling with noise-free data
Neural scaling is an effective method of measuring parameter efficiencies in neural networks. Here, different parameter sizes were investigated by changing the number of nodes in the hidden layers. Training and test losses were evaluated for data without added noise, to isolate the underlying expressive capabilities of each technique. Figure 4 reveals neural scaling at orders of 1.0 and 0.6 for the training and testing metrics of ChemKANs (where the order of neural scaling is defined as the power to which the loss decreases with respect to the number of parameters). While we might expect values up to 4 based on prior KAN [23] and KAN-ODE [29] studies, we remark here that the low-order neural scaling appears to indicate relatively saturated training of the ChemKAN with errors already around with just 72 parameters, rather than poor convergence which might otherwise have been indicated if low convergence rates were coupled with poor loss metrics.
@@
To further probe the nuances of ChemKAN’s convergence efficiency, we compare its results with those of DeepONet. The ChemKAN iterates roughly an order of magnitude slower but converges in fewer epochs (as originally discussed in [29]). Thus, to facilitate fair comparison the ChemKAN was trained only for 5,000 epochs, while the DeepONet was trained for 50,000 epochs. Two key distinctions between ChemKAN and DeepONet are as follows. First, as discussed earlier, the extremely sparse ChemKANs (toward the left half of Fig. 4(A)) are able to reach remarkably low error even with just 78 parameters, while the DeepONets see significantly worse performance at sparse parameterizations (also seen in the left half of Fig. 4(A)). Secondly, we note that the DeepONet sees significantly higher order neural convergence in the training results, allowing it to surpass the training performance of the ChemKAN at above 200 parameters, with remarkably strong training accuracy for the largest, 456-parameter DeepONet studied here. Linear fits with slopes are shown in Fig. 4(A) to illustrate this point, where the last ChemKAN and last two DeepONet points are excluded as they begin to plateau in training loss. From this observation, a large-enough DeepONet seems to outperform ChemKAN when looking only at the training losses.
@@
To further contextualize the structural efficiency, we discuss results for the testing error of the same neural convergence runs in Fig. 4(B). Here, we notice a significant departure from the training convergence rates in Fig. 4(A). Unlike the neural convergence for training data, the ChemKAN is seen to outperform the DeepONet at all sizes, with the largest ChemKAN leveling out and retaining nearly the same testing performance as the second-largest ChemKAN. The DeepONet, while enjoying a faster neural convergence rate below 308 parameters, notably fails to plateau and instead appears to diverge at higher parameter counts. When compared against the training results in Fig. 4(A), we observe two distinct modes of training saturation. Saturation, or the point where the linear fit no longer holds, appears to occur for the two largest DeepONets. For the ChemKAN, we might either interpret the entire profile to be saturated, or highlight the single largest network as the saturation point. Regardless, what we observe in these high-parameter networks is high robustness in the ChemKAN to overfitting (with a flat testing loss plateau), compared to the significant overfitting and divergence seen in the DeepONet past 308 parameters (i.e., the last two testing points seeing increasing loss).
@@
It is unsurprising that a standard deep learning technique begins to overfit a small dataset when given a large number of parameters. The DeepONet overfitting past saturation leads to further decreases in training loss accompanied by significant increases in testing loss. What we do find surprising, however, is ChemKAN’s apparent resilience to overfitting, even with similar parameter counts. In this context, its original failure to reach the same training performance as the DeepONet appears to be a strength of the method rather than a drawback, as it reaches a minimum value for both training and testing and then remains robust to superfluously added parameters, while the DeepONet clearly requires additional care to avoid overfitting. This pilot neural convergence test suggests that ChemKANs are robust to overfitting, and motivates further study of their capability in a second, more realistic model inference scenario.
III.1.3 Benefits of latent-dynamics modeling: noisy inference comparison
@@
To gain a deeper understanding of how ChemKAN outperforms DeepONet, we further test the training behaviors of ChemKAN and DeepONet with noisy data. Early stopping is carried out for both network structures at 10,000 epochs, to enable fair comparison and limit overfitting in the DeepONet cases. In all cases, a DeepONet with 308 parameters was used to compare against a ChemKAN with 156 parameters. The 308-parameter DeepONet was chosen based on the results of Fig. 4, where this was seen to be the largest DeepONet before the test loss began overfitting. The 156-parameter ChemKAN, meanwhile, was chosen to roughly match the DeepONet’s training and testing performance at zero noise. In other words, we chose the best-performing DeepONet size possible based on the preliminary neural convergence study, and then sized the ChemKAN according to the zero-noise performance of both networks. In more detail, the DeepONet had a three-layer branch network with eight nodes per layer and a two-layer trunk network with seven nodes in the first layer and eight in the second layer, with a final output layer converting these two eight-dimensional layers to the six-dimensional solution vector. The ChemKAN, meanwhile, had a single hidden layer with four nodes, two of which included multiplication operators () as per the standard LeanKAN formulation [35], and three gridpoints per activation.
@@
Figure 5(A) shows average training results after epochs across the 20 training and 10 testing cases at varied noise levels from to . As expected from Sec. III.1.2, the training MSE with 0% noise shows that the 308-parameter DeepONet slightly beats the 156-parameter ChemKAN in training performance, and is slightly worse in terms of reconstructing the unseen testing data.
@@
As increasing noise is added to the system, we see in the standard training and testing metrics of Fig. 5(A) that both networks unsurprisingly see increases in training and testing errors. Roughly, the increase in MSE error in the training losses scales with the square of the noise, as follows from Eq. 18. For example, the ChemKAN sees an increase in MSE between 0% noise and 1% noise of , where the second degree scaling of the MSE suggests that a 25 larger increase of might be expected between 0% noise and 5% noise. This is indeed observed, with an increase in this latter case of . Thus, the increase in training error for both networks as noise is added can be attributed to the effect of the noise itself on the loss function (Eq. 18), and does not appear to indicate any problems with the two networks’ capabilities to fit the increasingly noisy training data. For a direct comparison, The ChemKAN retains lower testing error throughout all tested noise values, and at noise and above, it is actually able to reach a lower training error than the DeepOnet. Looking at the big picture, however, results in these two metrics remain within a factor of two of each other at all noise levels, indicating largely similar performance.
@@
Upon further evaluation, we found that the training and testing MSEs evaluated on noisy datasets do not fully capture the effects of noise on useful model predictions, as overfitting can occur not only to the training conditions but also to the noise present in the data. To more effectively compare these frameworks, we introduce a noise-free MSE metric as was previously studied in the context of KANs [32],
@@
which differs from the test MSE in Eq. 18 in its use of the true, noise-free data rather than the noisy observations. This noise-free metric serves to quantify the capability of the two modeling approaches to accurately extract the true underlying model from noisy data, rather than overfit the noise or otherwise fail to deliver a useful model. With this metric, much more significant performance shifts can be seen in the noise-free testing values in Fig. 5(A). The impact of the added noise on the noise-free testing error for the ChemKAN is relatively small (a roughly increase from to noise), suggesting that the ChemKAN is able to extract the true underlying behavior well from the noisy datasets. We reiterate the significance of this result in the context of prior work [32], where standard KANs were demonstrated to fail when faced with small added noise. In contrast, the DeepONet sees a 5 increase in noise-free testing MSE from to noise, with its final noise-free testing loss 4.4 larger than that of the ChemKAN.
@@
The loss profiles in Figs. 5(B) provide further insight on the training dynamics that lead to this significant discrepancy in noise-free testing reconstructions. In the noise training cycle of Fig. 5(B), we see fairly standard behavior across the training and noise-free testing traces for the DeepONet and ChemKAN, with all quantities steadily decreasing for the entire duration of training. This is the expected result, as we have sized the DeepONet based on Fig. 4 to avoid any overfitting in the noise-free case. As we increase the amount of noise to , the training loss values are heavily penalized (due to the noisy data used in the computation of Eq. 18), while the noise-free testing values are slightly penalized but remain comparatively strong, indicating that both approaches are at least to a certain extent able to extract the true underlying model from the noisy data. While the training loss continues to drop quickly and then plateau in all four subplots, we see in the 7% noise case of Fig. 5(B) that the DeepONet noise-free testing loss dynamics begin to suffer, with a minimum value near 5,000 epochs and a slight upward trend toward later epochs, likely due to overfitting. In the 15% noise case of Fig. 5(B), this issue is further exacerbated, with an early minimum near 1,000 epochs followed by significant overfitting to the noisy data for the remainder of the training profile. The ChemKAN in both cases remarkably continues to drop its noise-free testing loss even while the DeepONet is overfitting, with late-epoch dynamics showing plateaued minimum values rather than the overfitting seen in the DeepONet. This echoes the behavior seen in Fig. 4, where the ChemKAN did not overfit and instead simply plateaued at its minimum training and testing errors.
@@
Reconstructed training data profiles are shown for the 15% noise case in Fig. 6. While the DeepONet and ChemKAN are both roughly able to find the unseen, noise-masked ground truth profile, a close inspection reveals not only better ChemKAN fits but also notably jagged profiles from the DeepONet as it attempts to overfit the noise present in the training data. These results suggest that the dynamical system exploitation inherent to the ODE-based framework of ChemKANs (and KAN-ODEs in general) can mitigate or even entirely resolve previously observed issues [32] regarding noisy data with KANs, and help to surpass the performance of standard DeepONets.
@@
In summary, we have demonstrated ChemKANs as a promising tool for model discovery in temperature-dependent chemical kinetic systems, especially with realistically noisy datasets. They show promise not only on the scientific side of the problem, where they were demonstrated here to have significant capability compared to a standard tool in discovering models hidden under noisy data, but also on the machine learning side of the problem, where we have demonstrated that the neural ODE implementation of KAN-ODEs and ChemKANs helps them to overcome the noisy data limitation recently shown in vanilla KANs [32].
III.2 Hydrogen combustion acceleration from homogeneous reactor data
@@
In our second case study, we investigate the use of ChemKANs as a reduced-order solver acceleration framework for known chemical models. We hypothesize based on previous high-order neural scaling results [29, 23] as well as the strong performance at low parameter counts seen in Fig. 4 that the use of a KAN structure instead of an MLP will allow for similar accuracy in network predictions with fewer parameters and lower cost. We additionally aim to evaluate the novel network architecture of the ChemKAN, and whether its physics-based, single-network structure is able to model behavior that required MLPs in the Neural ODE framework [17]. To reduce computational cost to the furthest extent possible, we use a single hidden layer of just three nodes for the core kinetic network (Eq. 13). This three-node hidden layer works as a latent representation of the system’s dynamics, compressing the information from 10 thermochemical states in the input and 29 reactions in the kinetic model. Leveraging our knowledge of the functional form of the governing equations (Eqs. 1 and 2) and their strong multiplicative behavior, we use here, defining all three hidden nodes using the multiplication operator.
@@
We begin in Fig. 7 with a demonstration of the ChemKAN homogeneous reactor reconstructions for one training case (left column, and K) and the unseen testing case (right column, and K), after the two-stage training process outlined in Sec. II.4.2. These reconstructions were generated entirely by the ChemKAN, given only the initial conditions. Overall the learned model successfully predicted the temperature and mass fractions in both cases, with no notable deterioration in the testing case (as expected from the strong testing results and robustness to overfitting observed in the previous biodiesel investigation). We further emphasize that the ChemKAN was largely able to capture the behavior of the low-concentration and highly reactive species , , and (Fig. 7(C, F)) that were neglected in ChemNODE [17].
@@
A broader comparison is shown in Fig. 8(A), where the MSE (in the normalized u units) is plotted across not only the initial set of 35 training and 1 testing initial conditions, but a wider set of 441 total initial conditions (406 of which were unseen during training) at a finer resolution in the same range. The low-temperature initial conditions see near-perfect reconstructions, as the ignition delays there are larger than the studied time window, leading to smooth gradients and easily trainable, near-isothermal behavior. In the remainder of the domain, strong performance is seen at all training conditions, and additionally at the single testing condition plotted in Fig. 7. In terms of generalization to intermediate temperatures and equivalence ratios, the ChemKAN performed very well throughout the vast majority of the domain, with many testing points even surpassing the accuracy of nearby training points at and above 1050 K. We notice that toward the slower-igniting cases, however, the ChemKAN struggled more with generalization. At 1000 K, for example, all six training points saw strong MSE values in the range, although a few of the intermediate equivalence ratios suffered. At 987.5 K, however (one tick below 1000 K), all testing points saw poorer performance in the range. We can conclude from these results that the ChemKAN retains its strong capability to generalize in the more challenging H2-air combustion case (as was originally reported in the biodiesel modeling case), but with practical limits in the colder, more temperature-sensitive ignition cases. While the ChemKAN’s ability to accurately learn the six studied training points at 1000 K suggests that it is capable of tracking ignition behavior through the cooler regions, its relatively poor performance elsewhere in the initiation-sensitive regime suggests that a non-uniform training grid with denser sampling toward such cooler regions is needed to fully resolve this behavior and provide more accurate results when applied in combustion CFD simulations that require accurate ignition behavior.
@@
We finally plot the actual and KAN-ODE predicted ignition delays in Fig. 8(B), for the 30 studied cases that ignited (the lowest temperature cases did not see ignition given the time span of 0.6 ms). Ignition here is defined as the point of maximum temperature rise rate [17]. Accuracy is strong across the board, even in the testing case.
@@
This collection of results shows that the ChemKAN structure was able to accurately learn the dynamics of all nine species and temperature scalars across the same set of initial reactor conditions as was studied using traditional MLP-based Neural ODEs in ChemNODE [17]. Compared to the six species plus temperature scalars learned there via seven unique MLP networks with 91 parameters each (according to standard MLP parameterizations, 637 total parameters), the current ChemKAN was able to learn the complete set of thermochemical scalars (nine species plus temperature) using a single, 344-parameter network. While training took place in two stages to decouple the kinetic and thermodynamic behavior and facilitate convergence, the final network remains a single cohesive structure with shared information across all nodes, eliminating the redundancies in repeated yet isolated 91-parameter MLPs.
@@
Finally, regarding computational efficiency, we report that the average time to solve all 36 homogeneous reactor conditions in the Arrhenius.jl combustion solver package [52] was 2 faster when switching from the detailed chemistry to the reduced ChemKAN framework. The total number of time steps in the integration process remains largely unchanged, as the ChemKAN is solving the same full-dimensional thermochemical state. The 2 speedup is predominantly achieved through faster gradient computation (i.e., each time step is faster to compute), facilitated by the ChemKAN’s compression of 29 reactions into just three hidden nodes and a handful of sparse activations. By learning the relationship between the current thermochemical state and the chemical source terms, the ChemKAN is capable not only of predicting ignition delay times and homogeneous reactor solution profiles 2 faster than the detailed model, but also of generalizing to other simulation conditions when coupled to flow solvers, including simple laminar flames and complex 2-D and 3-D turbulent combustion conditions. Such downstream uses of similar surrogate machine learning models were discussed and tested in previously [16, 17], where a 2 speedup in the chemical solver (which is often the most computationally expensive component in a reacting flow simulation) implies the potential for substantial acceleration unlocked by ChemKANs while retaining the full-sized, detailed solution state vector. While slightly slower than the 2.3 speedup reported in ChemNODE [17], we reiterate that the ChemKAN solves for an additional three minor species (including the key H radical). A summarized comparison of ChemKAN and ChemNODE is provided in Table 1.
# of nets | # of params | Species modeled | Speed-up vs. true model | |
ChemNODE[17] | 7 | 637 | H2, O2, H2O, N2, O, OH | 2.3 |
ChemKAN (our work) | 1 | 344 | H2, O2, H2O, N2, O, OH, H, HO2, H2O2 | 2.0 |
III.3 Current limitations and areas for further research